### v1

@@ 2.6 Combination of similar impactors
On the last stage we check if some of the found VIs are essentially the same and point to the same possible collision. For this we integrate the orbits of each VI until the impact, saving their positions every 30 days. If at each integration step several VIs are close to each other (distance $10^{-3}$ au), we take only the one with the highest impact probability value. We publish a possible collision if the impact probability is higher than $10^{-10}$. The reason why this threshold is significantly higher than $5\cdot 10^{-8}$ threshold for the maximal impact probability is that if the maximal impact probability is too small it means the uncertainty of the asteroid around this area is too large. On the other hand, if the maximal impact probability is not too small but the actual impact probability of a VI is very low, this means that the uncertainty region only grazes the Earth rather than passing through its center. In such cases, even small changes in the nominal orbital solution can lead to large fluctuations in the estimated impact probability, which is why these cases are interesting.
@@ 3.1 Classical approach
$q=\frac{b}{\sqrt{1+\frac{v_{s}^{2}}{v_{rel}^{2}}}},$ (11)
where $v_{s}\approx 11.186\ \mathrm{km\ s^{-1}}$ is the Earth’s surface escape velocity and $b$ is the impact parameter (the distance on the target plane between the projections of the Earth’s and asteroid’s centers).
@@ 3.2 Beyond classical approach
To address this, we adopted a modified approach similar to Milani et al. (2005a). We extend the integration until the epoch of closest approach and construct a ”modified target plane (MTP)” perpendicular to the relative velocity at that instant. If the asteroid collides with the Earth, the MTP is defined at the impact point. Using the relative position and velocity at closest approach, we compute the unperturbed velocity $v_{\infty}$ and derive the corresponding unperturbed target-plane coordinates $(\xi_{TP^{*}},\zeta_{TP^{*}})$ from the MTP coordinates $(\xi_{MTP},\zeta_{MTP})$:
@@
3.2.1 Search for virtual impactor
To confirm whether a potential encounter corresponds to a genuine collision, we follow the method of Milani et al. (2005a). The 6D state vector at $t$ relative to the nominal state vector of the asteroid can be partitioned into $\bm{r},\bm{s}]$, where $\bm{r}=(\xi_{TP},\zeta_{TP})$ are target-plane coordinates and $\bm{s}$ is the complementary 4D vector.
@@ 3.2.1 Search for virtual impactor
According to (Milani et al., 2005a) , vector $\bm{s_{*}}$ can be found from
@@
where $\mathbf{W}$ is the $6\times 6$ rotation matrix into the TP2 frame. As in Section 2.5 we also consider $\frac{3}{4}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$ and $\frac{1}{2}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$ and select the closest one to the Earth at time $t$. If all these three are more distant than $\bm{x}_{0*}$ at time $t$ we take $\frac{1}{6}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$.
@@ 4 Practical realization and discussion
In this version of the system we used Everhart (1974) integrator of 15-th order. In the model we included gravitational influence of 8 major planets (and Pluto for historical reasons). We also considered the Earth and the Moon separately. The orbits of asteroids and their covariance matrices are taken from DynAstVO database [^3] https://epn.imcce.fr/dynastvo.html Desmars et al. (2017).
@@ 4.1 Completeness problem
The completeness of our system is determined by the threshold of impact probability that can be detected certainly. Following (Milani et al., 2005a) one can estimate the generic completion level: the highest impact probability that can be missed by our monitoring system if the uncertainty trail on the target plane is fully linear.
Under these assumptions, a possible collision can be missed if both representatives found in Section 2.4 are further than 1 au from the Earth. This literally means that the distance between two subsequent VAs at the time of this possible collision exceeds 2 au along the orbit. According to the sampling of VAs along the LOV, the maximal integral probability of a Gaussian distribution on the interval between two subsequent VAs is approximately $3\times 10^{-4}$. This is slightly higher than if we sampled VAs along the LOV according to Gaussian distribution. In this case, the probability would be $2\cdot 10^{-4}$. Our objective is hence the following: If we took a sample of 5001 VAs according to the Gaussian distribution, the latest VAs would be only 3.54 sigmas away from the nominal state vector ($l_{2500}=3.54$ from equation 2). By sampling the Laplace distribution $l_{2500}=5.53$ instead, we have VAs much further from the nominal state vector, without loosing much in terms of completeness. Hence, the highest impact probability that can be missed is approximately:
@@ 4.2 Comparison to Sentry
To evaluate the robustness of our system, we performed detailed comparisons using five representative asteroids: 2005 QK76, 2008 JL3, 2023 DO, 2025 JU, and 2000 SG344. We selected these asteroids from the web-page Sentry of Central for Near-Earth Objects Study (CNEOS) [^4] https://cneos.jpl.nasa.gov/sentry. We were looking for objects at the top of the list with high cumulative impact probability and significant number of possible impacts with probabilities higher than $5\times 10^{-7}$ in order to be able to make a comparison. The asteroids have different observational arcs (from 1.7 days to 1.4 years), and different number of observations (from 14 to 52). For each case, we adopted the same orbital solutions as those used by NASA’s Sentry-II system, retrieved via the NASA SSD API [^5] https://ssd-api.jpl.nasa.gov on September 9, 2025, so that the results would be directly comparable.
@@
In our approach, $\sigma_{VI}$ quantifies how far a VI lies from the nominal orbit in terms of standard deviations and is defined as
@@
where $\bm{x}_{0*}$ is the state vector of the VI, $\bm{x}_{0}$ is the nominal state vector, and $\mathbf{C}_{0}$ is the covariance matrix. Sentry-II estimates $\sigma{VI}$ using the root-mean-square residuals of the observational fit for the impacting solution.
@@
Table 1: Comparison table for asteroid 2025 JU.
@@
2074-10-2.3 4.10e-05 1.091 4.30e-05 1.539
2074-10-2.3 1.49e-05 0.998 1.50e-05 1.500
2076-10-1.9 1.71e-07 3.268 1.70e-07 3.855
2079-10-2.5 1.91e-08 0.949 2.00e-08 1.362
2079-10-2.5 8.04e-08 0.947 8.00e-08 1.193
2079-10-2.5 6.25e-07 0.742 6.80e-07 0.956
2079-10-2.5 5.68e-07 0.741 5.50e-07 0.937
2091-10-2.4 5.50e-10 1.182
2091-10-2.5 3.56e-09 1.830
2103-10-3.3 5.25e-09 3.011
2108-10-2.6 2.21e-08 1.164 2.90e-08 1.929
2108-10-2.5 7.23e-08 1.170 6.70e-08 1.471
2108-10-2.6 3.13e-07 1.175 3.10e-07 1.675
4.2.1 Asteroid 2025 JU
Asteroid 2025 JU is newly discovered, with 38 observations spanned over 35 days. Its cumulative impact probability is about $6\times 10^{-5}$, not very high and it also has only a few significant possible impacts. Table 1 shows an iconic agreement between two systems with Sentry loosing three possible collisions with a probability less than $5.3\times 10^{-9}$ (way below the mentioned completeness level).
@@ 4.2.2 Asteroid 2005 QK76
Asteroid 2005 QK76 has been discovered on August 30 2025. It has only 14 observations spanned over 1.67 days. Such a short observation arc makes its orbit highly uncertain and a good test for comparison. Sentry-II does not use the covariance matrix only the observations itself, while NEOForCE uses the covariance matrix in coordinates and velocities at the mean epoch of observations. As one can see from Table 2, there is no significant difference in the impact probabilities of the possible collisions found. However, the values of $\sigma_{VI}$ are substantially different, with larger values for Sentry-II. Sentry-II did not find the possible collision on February 26, 2038 with impact probability of $1.7\times 10^{-7}$, but this is at the completeness limit of Sentry-II.
@@ 4.2.4 Asteroid 2008 JL3
This asteroid was observed over a short arc of only 7 days, yielding 35 measurements in total. Discovered on May 2, 2008, it has 44 possible collision solutions listed in Sentry-II. A comparison of our system with Sentry-II (Table NEOForCE: Near-Earth Objects’ Forecast of Collisional Events) shows that NEOForCE identified 87 possible collisions but did not recover 3 mentioned by Sentry-II, all with impact probabilities below $4.6\times 10^{-9}$ — way below our completeness threshold. In contrast, Sentry-II missed 46 possible collisions, including one with the highest reported impact probability of $1.26\times 10^{-6}$ on April 28, 2109, as well as three others with probabilities exceeding $10^{-7}$.
@@ 4.2.5 Asteroid 2000 SG344
Asteroid 2000 SG344 turns out to be a hard test for our systems. It was discovered on 29 September 2000 by the LINEAR survey at Socorro, New Mexico. The orbital solution used today by NASA JPL is based on observations spanning from 15 May 1999 to 3 October 2000, giving it an observational arc of about 1.4 years, with only a modest number of measurements available for orbit fitting (31 observations). Its orbital period is nearly one year, the inclination relative to the ecliptic is almost zero, and the aphelion lies just beyond Earth’s orbit. These parameters result in very low encounter velocities with Earth (about 1.4 km/s) and a continuum of predicted close approaches after 2070. Because of this unusual dynamical configuration, 2000 SG344 has been the subject of a dedicated impact-hazard study (Chodas & Chesley, 2001).

### v2

@@ 2.6 Combination of similar impactors
On the last stage we check if some of the found VIs are essentially the same and point to the same possible collision. For this we integrate the orbits of each VI until the impact, saving their positions every 30 days. If at each integration step several VIs are close to each other (distance $10^{-3}$ au), we take only the one with the highest impact probability value. We publish a possible collision if the impact probability is higher than $10^{-10}$. The reason why this threshold is significantly higher than $5\times 10^{-8}$ threshold for the maximal impact probability is that if the maximal impact probability is too small it means the uncertainty of the asteroid around this area is too large. On the other hand, if the maximal impact probability is not too small but the actual impact probability of a VI is very low, this means that the uncertainty region only grazes the Earth rather than passing through its center. In such cases, even small changes in the nominal orbital solution can lead to large fluctuations in the estimated impact probability, which is why these cases are interesting.
2.7 Short summary of the method
Below we provide a concise overview of the NEOForCE impact-monitoring algorithm. A schematic illustration of the work is shown in Fig. 4.
  1. 1.
    Sampling. We create many virtual asteroids (VAs) along the Line of Variations (LOV) at the mean observation epoch (Section 2.1).
  2. 2.
    Initial screening. We make a first quick test to see whether the asteroid can ever come close to the Earth. For this, we compute the minimum MOID for a small group of VAs (Section 2.2).
  3. 3.
    Propagation and close-approach detection. We integrate the orbit of each VA into the future. We record every moment when the Earth comes close to the osculating orbit of a VA (osculating ellipse), together with the time and distance (Section 2.3).
  4. 4.
    Curvilinear-region analysis. For every case where the distance is smaller than 0.05 au, we build the curvilinear uncertainty region of the VA and compute its distance to the Earth. We also estimate the maximum impact probability that this VA could have (Section 2.3).
  5. 5.
    Selection of representative scenarios. From the previous step, we select only the most important situations: VAs that are neighbors along LOV and bracket the Earth at approximately the same time, have small distances between their uncertainty regions and the Earth, and show a noticeable impact probability (Section 2.4).
  6. 6.
    Search for virtual impactors. For each selected case, we try to improve the VA’s state vector so that the VA hits the Earth. We do this step by step, coming closer to an impact solution. When the VA enters the Earth’s sphere of influence at the probable collision time, we use the target-plane analysis (Sections 2.5 and Section 3).
  7. 7.
    Combine similar impactors. If two virtual impactors stay very close to each other on their way toward the Earth, we treat them as the same solution and keep the one with the higher impact probability (Section 2.6).
@@ 3.1 Classical approach
$q=\frac{b}{\sqrt{1+\frac{v_{s}^{2}}{v_{\infty}^{2}}}},$ (11)
where $v_{s}\approx 11.186\ \mathrm{km\ s^{-1}}$ is the Earth’s surface escape velocity, $v_{\inf}$ is the unperturbed geocentric velocity and $b$ is the impact parameter (the distance on the target plane between the projections of the Earth’s and asteroid’s centers).
@@ 3.2 Beyond classical approach
To address this, we adopted a modified approach similar to Milani et al. (2005b). We extend the integration until the epoch of closest approach and construct a ”modified target plane (MTP)” perpendicular to the relative velocity at that instant. If the asteroid collides with the Earth, the MTP is defined at the impact point. Using the relative position and velocity at closest approach, we compute the unperturbed velocity $v_{\infty}$ and derive the corresponding unperturbed target-plane coordinates $(\xi_{TP^{*}},\zeta_{TP^{*}})$ from the MTP coordinates $(\xi_{MTP},\zeta_{MTP})$:
@@
3.3 Search for virtual impactor
To confirm whether a potential encounter corresponds to a genuine collision, we follow the method of Milani et al. (2005b). The 6D state vector at $t$ relative to the nominal state vector of the asteroid can be partitioned into $\bm{r},\bm{s}]$, where $\bm{r}=(\xi_{TP},\zeta_{TP})$ are target-plane coordinates and $\bm{s}$ is the complementary 4D vector.
@@ 3.3 Search for virtual impactor
According to (Milani et al., 2005b) , vector $\bm{s_{*}}$ can be found from
@@
where $\mathbf{W}$ is the $6\times 6$ rotation matrix into the TP2 frame. As in Section 2.5 we also consider $\frac{3}{4}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$ and $\frac{1}{2}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$ and select the closest one to the Earth at time $t$. If all these three are more distant than $\bm{x}_{0*}$ at time $t$ we take $\frac{1}{10}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$.
@@ 4 Practical realization and discussion
In this version of the system we used collocational integrator Lobbie of 14-th order (Avdyushev, 2022) for integrating the orbits. In the model we included gravitational influence of 8 major planets (and dwarf planet Pluto, for historical reasons), as well as post-Newtonian formalism for the Sun only. We also considered the Earth and the Moon separately. The orbits of asteroids and their covariance matrices are taken from DynAstVO database [^4] https://epn.imcce.fr/dynastvo.html Desmars et al. (2017) for the automated impact monitoring.
@@ 4.1 Completeness problem
The completeness of our system is determined by the threshold of impact probability that can be detected certainly. Following (Milani et al., 2005b) one can estimate the generic completion level: the highest impact probability that can be missed by our monitoring system if the uncertainty trail on the target plane is fully linear.
Under these assumptions, a possible collision can be missed if both representatives found in Section 2.4 are further than 1 au from the Earth. This literally means that the distance between two subsequent VAs at the time of this possible collision exceeds 2 au along the orbit. According to the sampling of VAs along the LOV, the maximal integral probability of a Gaussian distribution on the interval between two subsequent VAs is approximately $3\times 10^{-4}$. This is slightly higher than if we sampled VAs along the LOV according to Gaussian distribution. In this case, the probability would be $2\times 10^{-4}$. Our objective is hence the following: If we took a sample of 5001 VAs according to the Gaussian distribution, the latest VAs would be only 3.54 sigmas away from the nominal state vector ($l_{2500}=3.54$ from equation 2). By sampling the Laplace distribution $l_{2500}=5.53$ instead, we have VAs much further from the nominal state vector, without loosing much in terms of completeness. Hence, the highest impact probability that can be missed is approximately:
@@ 4.2 Comparison to Sentry
To evaluate the robustness of our system, we performed detailed comparisons using five representative asteroids: 2005 QK76, 2008 JL3, 2023 DO, 2008 EX5, and 2000 SG344. We selected these asteroids from the web-page Sentry of Central for Near-Earth Objects Study (CNEOS) [^5] https://cneos.jpl.nasa.gov/sentry. We were looking for objects at the top of the list with high cumulative impact probability and significant number of possible impacts with probabilities higher than $5\times 10^{-7}$ in order to be able to make a comparison. The asteroids have different observational arcs (from 1.7 days to 1.4 years), and different number of observations (from 14 to 52). For each case, we adopted the same orbital solutions and the same covariance matrices as those used by NASA’s Sentry-II system, retrieved via the NASA SSD API [^6] https://ssd-api.jpl.nasa.gov on September 9, 2025, so that the results would be directly comparable.
@@
$\sigma_{VI}$ quantifies how far a VI lies from the nominal orbit in terms of standard deviations and is defined as
@@
where $\bm{x}_{0*}$ is the state vector of the VI, $\bm{x}_{0}$ is the nominal state vector, and $\mathbf{C}_{0}$ is the covariance matrix.
@@
Table 1: Comparison table for asteroid 2008 EX5.
@@
2056-10-8.9 7.04e-08 0.504 7.00e-08 1.028
2059-10-9.5 5.38e-08 0.505
2061-10-9.0 3.00e-08 0.501 3.00e-08 1.400
2062-10-9.5 6.64e-07 0.562 7.80e-07 0.604
2063-10-9.5 5.09e-08 0.498 4.80e-08 0.618
2065-10-9.0 4.41e-08 0.506 4.70e-09 0.564
2065-10-09.09 4.50e-08 1.085
2067-10-9.6 4.11e-08 0.503
2067-10-9.8 9.43e-08 0.710
2068-10-9.0 5.09e-07 0.551 6.90e-07 0.853
2068-10-08.85 3.50e-10 2.175
2069-10-9.1 1.02e-07 0.255 1.00e-07 1.324
2070-10-09.39 6.20e-08 0.649
2071-10-9.7 2.10e-06 0.371 2.00e-06 0.644
2071-10-09.63 3.40e-08 1.238
2072-10-9.0 1.94e-05 0.919 2.70e-05 0.922
2072-10-8.9 1.37e-06 1.533 1.40e-06 1.685
2074-10-9.5 4.62e-07 0.481 4.80e-07 0.592
2074-10-9.4 1.26e-07 1.615 1.30e-07 1.707
2075-10-9.8 2.19e-08 0.917 6.70e-10 1.449
2076-10-8.9 6.60e-08 0.394 6.30e-08 0.447
2076-10-8.9 2.52e-08 0.420
2077-10-9.1 3.11e-07 0.542 3.50e-07 0.699
2079-10-9.8 3.74e-08 0.266 3.80e-07 0.286
2079-10-9.6 3.45e-08 0.578 3.40e-08 1.161
2080-10-9.0 5.28e-08 0.715
2081-10-9.1 8.56e-08 1.579
2082-10-9.3 9.08e-08 1.620 9.30e-08 2.029
2082-10-9.3 1.80e-10 0.576
2083-10-9.7 1.89e-05 0.844 2.00e-05 1.002
2083-10-09.66 1.30e-08 0.617
2084-10-8.9 5.56e-08 1.582 5.40e-08 1.619
2086-10-9.5 4.42e-10 0.257
2088-10-09.02 5.40e-09 1.445
2090-10-9.5 5.44e-08 0.030 5.80e-08 1.282
2093-10-09.25 2.30e-09 0.045
4.2.1 Asteroid 2008 EX5
Asteroid 2008 EX5 is newly discovered, with 61 observations spanned over 32 days. Its cumulative impact probability is about $4.5\times 10^{-5}$, not very high and it also has only a few significant possible impacts. Table 1 shows an agreement between two systems with Sentry loosing 8 possible collisions with a probability less than $1\times 10^{-7}$ (below the mentioned completeness level) and NEOForCE looses 7 possible collisions with probabilities less than $5\times 10^{-8}$, below the threshold for maximal impact probability.
@@ 4.2.2 Asteroid 2005 QK76
Asteroid 2005 QK76 has been discovered on August 30 2025. It has only 14 observations spanned over 1.67 days. Such a short observation arc makes its orbit highly uncertain and a good test for comparison. As one can see from Table 2, there is no significant difference in the impact probabilities of the possible collisions found. However, the values of $\sigma_{VI}$ are substantially different, with larger values for Sentry-II. Sentry-II did not find the possible collision on February 26, 2038 with impact probability of $1.7\times 10^{-7}$, but this is at the completeness limit of Sentry-II.
@@ 4.2.4 Asteroid 2008 JL3
This asteroid was observed over a short arc of only 7 days, yielding 35 measurements in total. Discovered on May 2, 2008, it has 44 possible collision solutions listed in Sentry-II. A comparison of our system with Sentry-II (Table 4.2.5) shows that NEOForCE identified 87 possible collisions but did not recover 3 mentioned by Sentry-II, all with impact probabilities below $4.6\times 10^{-9}$ — way below our completeness threshold. In contrast, Sentry-II missed 46 possible collisions, including one with the highest reported impact probability of $1.26\times 10^{-6}$ on April 28, 2109, as well as three others with probabilities exceeding $10^{-7}$.
@@ 4.2.5 Asteroid 2000 SG344
Asteroid 2000 SG344 turns out to be a hard test for our systems. It was discovered on 29 September 2000 by the LINEAR survey at Socorro, New Mexico. The orbital solution used today by NASA JPL is based on observations spanning from 15 May 1999 to 3 October 2000, giving it an observational arc of about 1.4 years, with only a modest number of measurements available for orbit fitting (31 observations). Its orbital period is nearly one year, the inclination relative to the ecliptic is almost zero, and the aphelion lies just beyond Earth’s orbit. These parameters result in very low encounter velocities with Earth (about 1.4 km/s) and a continuum of predicted close approaches after several resonance returns after 2070. Because of this unusual dynamical configuration, 2000 SG344 has been the subject of a dedicated impact-hazard study (Chodas and Chesley, 2001).
@@
For the reasons discussed above, the computation time for asteroid 2000 SG344 was significantly higher (more then 6x longer than for 2008 JL3). Table 6 reports the total computation time, obtained in parallel mode using two threads on an Intel(R) Xeon(R) Gold 6230 CPU at 2.10 GHz, together with the number of situations in which the virtual impactor (VI) search procedure was triggered.
Table 5: Comparison table for asteroid 2008 JL3
Date $P_{NEOForCE}$ $\sigma_{VI}$ $P_{NASA}$ $\sigma_{VI_{NASA}}$
2027-04-31.3 1.39e-04 0.111 1.50e-04 0.120
2030-04-30.8 3.00e-08 4.242 3.00e-08 4.286
2034-04-30.5 1.67e-07 0.105 1.60e-07 0.110
2043-05-01.05 4.60e-09 0.117
2047-05-01.52 3.20e-09 0.116
2048-04-30.5 1.32e-08 0.115
2051-04-31.4 7.34e-08 0.104
2061-04-31.1 5.74e-09 0.115
2063-05-2.0 2.28e-07 0.092 2.10e-07 0.096
2065-04-27.7 8.02e-08 0.911 7.90e-08 0.920
2071-04-30.3 2.63e-06 1.022 2.60e-06 1.032
2072-04-30.5 2.03e-08 0.117
2075-05-2.1 8.78e-08 1.598 1.00e-07 1.614
2075-05-2.0 9.28e-08 0.081
2077-04-28.5 9.27e-09 2.222
2080-04-26.93 2.40e-09 0.193
2081-04-31.5 1.39e-08 0.013
2082-04-27.4 4.08e-09 0.084 4.60e-09 0.153
2083-05-1.8 1.60e-08 0.095
2087-04-31.2 2.03e-08 2.145
2090-05-2.2 7.96e-08 0.099 7.70e-08 0.098
2090-05-1.6 1.25e-09 0.158 5.80e-09 0.167
2090-05-1.6 3.11e-07 0.193 2.80e-07 0.195
2090-04-31.5 5.40e-07 0.195 5.50e-07 0.197
2090-04-31.2 6.47e-07 0.229 5.80e-07 0.230
2090-04-31.5 5.52e-07 1.958 5.30e-07 1.978
2090-04-31.1 2.27e-08 0.188
2090-05-1.5 1.14e-08 2.291
2090-05-1.8 7.40e-08 2.792
2090-05-1.9 4.76e-08 3.011
2091-04-28.8 9.86e-09 0.919 9.10e-09 0.923
2092-04-28.9 2.72e-08 1.041 3.10e-08 1.052
2092-04-28.5 2.75e-08 0.189
2093-05-1.6 1.32e-08 0.166
2095-04-30.8 1.71e-08 0.085 1.60e-08 0.092
2095-04-31.4 3.21e-08 0.139 2.70e-08 0.139
2098-04-31.0 1.45e-08 0.190
2099-04-31.0 6.33e-09 2.141
2099-04-30.8 1.53e-08 0.084
2099-04-30.9 1.53e-08 1.023
2100-04-27.8 1.87e-07 1.267 1.80e-07 1.283
2100-04-27.9 2.63e-06 0.264 2.50e-06 0.275
2100-04-27.9 8.81e-07 0.174 8.60e-07 0.176
2100-04-27.8 1.25e-07 0.131 1.10e-07 0.134
2100-04-27.4 9.47e-10 2.651
2100-04-27.8 1.08e-08 2.297
2102-05-1.8 1.51e-08 0.231 3.40e-09 2.267
2103-04-28.3 4.90e-08 2.214
2103-04-28.5 1.86e-09 0.096
2109-04-27.9 2.55e-07 0.594 2.30e-07 0.601
2109-04-28.1 6.88e-08 1.089 6.50e-08 1.103
2109-04-28.1 4.91e-09 0.660
2109-04-28.0 1.26e-06 0.831
2109-04-28.2 8.90e-08 1.240
2110-05-2.9 4.12e-08 2.835 3.90e-08 2.864
2110-05-2.5 2.61e-08 2.055
2110-05-2.4 3.34e-08 1.925
2110-05-2.4 2.89e-08 1.463
2110-05-2.4 2.01e-09 0.190
2112-04-28.0 1.93e-07 0.210 1.80e-07 0.214
2112-04-27.9 3.18e-08 0.221 3.90e-10 0.149
2112-04-28.4 5.23e-07 0.906 4.90e-07 0.906
2112-04-28.3 2.49e-07 2.547 2.50e-07 2.573
2112-04-28.4 1.55e-07 0.903
2113-04-31.4 4.56e-07 0.115 4.50e-07 0.117
2113-05-1.8 4.37e-08 0.027 9.30e-09 0.054
2113-05-2.2 8.91e-10 3.718
2113-05-1.8 1.68e-07 2.495
2113-05-2.2 2.41e-08 0.005
2113-05-1.9 7.93e-09 0.147
2114-04-31.5 9.87e-09 0.084 9.50e-09 0.116
2114-05-1.6 1.54e-09 2.235
2115-04-28.5 3.32e-09 0.181
2115-04-28.4 1.60e-08 0.980
2116-04-30.1 1.39e-06 1.324 1.30e-06 1.338
2116-04-30.4 2.92e-07 0.025 2.60e-07 0.035
2117-05-2.2 1.85e-07 0.060 1.70e-07 0.061
2117-04-31.3 6.86e-10 0.188
2119-05-3.5 3.68e-08 0.234 3.50e-08 0.236
2119-05-3.5 9.36e-07 0.385 8.20e-07 0.389
2119-05-3.8 1.65e-10 0.106
2119-05-3.6 1.05e-07 2.012
2120-04-29.0 7.36e-10 2.812
2121-04-28.2 3.90e-08 0.178
2121-04-28.0 4.69e-09 0.192
2122-05-3.6 2.11e-08 0.091 2.90e-07 0.238
2122-05-3.1 2.18e-06 0.362 2.00e-06 0.366
2122-05-3.5 5.13e-07 2.030 5.10e-07 2.051
2123-04-30.7 9.69e-09 0.084
2124-04-27.9 1.47e-09 0.163
Table 6: Computation times and the number of VI searches for 5 considered asteroids.
Asteroid $N_{VI}$ $t_{comp}$, s
2008 EX5 965 2738.8
2005 QK76 555 2627.7
2023 DO 2118 6216.5
2008 JL3 984 2361.9
2000 SG344 8530 15524.7
’Asteroid’ is the asteroid designation, $N_{VI}$ is the number of situations in which the virtual impactor (VI) search procedure was triggered, $t_{comp}$ is the total computation time, obtained in parallel mode using two threads on an Intel(R) Xeon(R) Gold 6230 CPU at 2.10 GHz.
A summary of the comparison for all five considered asteroids with NASA is shown in Fig. 6. For asteroids 2008 EX5, 2005 QK76, 2023 DO, and 2008 JL3, we find good agreement between the two systems, with impact probability values consistent down to the $10^{-7}$ level, except for a few missing collision cases. For asteroid 2000 SG344, the agreement is noticeably worse and holds only down to about the $10^{-6}$ level. In this case, several potential collisions at the $10^{-5}$ probability level are missed by the NASA system. This behavior is likely caused by the very small relative velocity during close encounters, which reduces the completeness of the monitoring system (see Sect. 4.1).
Figure 6: Comparison between the NEOForCE impact monitoring system and Sentry-II.This figure summarizes the results from Tables 1–5. We show only impact probabilities larger than $10^{-9}$. A point located on the horizontal line at $P_{NASA}=10^{-9}$ indicates a potential impact identified by NEOForCE but not detected by Sentry-II. Conversely, a point on the vertical line indicates an impact detected by Sentry-II but not found by NEOForCE.